# Zero-Order Sharpness-Aware Minimization

### v1

@@ 6 Conclusion
In this paper, we propose ZOSA, a novel zero-order sharpness-aware minimization framework for efficient prompt tuning of large language models in resource-constrained environments. Our framework integrates batched Rademacher perturbations for gradient estimation, adaptive loss-variance scaling for stability, and sharpness-aware mechanisms to target flat minima. Theoretical analysis establishes $O(1/\sqrt{T})$ convergence under smoothness and bounded variance assumptions, with PAC-Bayesian bounds linking sharpness control to enhanced generalization. Empirical evaluations on synthetic high-dimensional functions and zero-order prompt fine-tuning across GLUE benchmarks validate ZOSA’s superiority, showing faster convergence, higher cosine similarity in gradient estimates, and enhanced accuracy/F1 scores compared to adaptive ZO baselines like ZO-AdaMM and evolutionary methods. These results underscore ZOSA’s robustness in noisy, high-dimensional landscapes, making it a practical solution for zero-order LLM adaptation.
@@ A.1 Fast Zeroth-Order Optimization
ZOSA approximates gradients using a finite difference method along random directions. For a parameter vector $\theta\in\mathbb{R}^{d}$ and a loss function $L(\theta)$, ZOSA estimates the gradient projection in $m$ random Rademacher directions $u_{i}$, where each component of $u_{i}$ is independently +1 or -1 with equal probability (i.e., $u_{i}$ is a Rademacher vector):
$\hat{g}_{i}=\frac{L(\theta+\epsilon u_{i})-L(\theta)}{\epsilon}u_{i}.$ (26)
This is a one-sided finite difference approximation, which reduces the number of forward passes compared to two-sided methods while maintaining effective gradient estimation. The expected value of this estimator, over the randomness of $u_{i}$, approximates the true gradient:
$\mathbb{E}[\hat{g}_{i}]=\mathbb{E}\left[\frac{L(\theta+\epsilon u_{i})-L(\theta)}{\epsilon}u_{i}\right].$ (27)
@@
Using a Taylor expansion around $\theta$:
$L(\theta+\epsilon u_{i})=L(\theta)+\epsilon\nabla L(\theta)^{\top}u_{i}+\frac{\epsilon^{2}}{2}u_{i}^{\top}H(\theta)u_{i}+O(\epsilon^{3}).$ (28)
Subtracting $L(\theta)$:
$\frac{L(\theta+\epsilon u_{i})-L(\theta)}{\epsilon}=\nabla L(\theta)^{\top}u_{i}+\frac{\epsilon}{2}u_{i}^{\top}H(\theta)u_{i}+O(\epsilon^{2}).$ (29)
Thus:
$\hat{g}_{i}=(\nabla L(\theta)^{\top}u_{i})u_{i}+\frac{\epsilon}{2}(u_{i}^{\top}H(\theta)u_{i})u_{i}+O(\epsilon^{2})u_{i}.$ (30)
Since $\mathbb{E}[u_{i}]=0$ and $\mathbb{E}[u_{i}u_{i}^{\top}]=I_{d}$ for Rademacher vectors, the expectation is:
$\mathbb{E}[\hat{g}_{i}]=\nabla L(\theta)+\frac{\epsilon}{2}\mathbb{E}[(u_{i}^{\top}H(\theta)u_{i})u_{i}]+O(\epsilon^{2}).$ (31)
The second term introduces a bias of order $O(\epsilon)$, as $\mathbb{E}[(u_{i}^{\top}Hu_{i})u_{i}]$ is generally non-zero but bounded for smooth Hessians $H$. Averaging over $m$ directions:
$\hat{g}=\frac{1}{m}\sum_{i=1}^{m}\hat{g}_{i},\quad\mathbb{E}[\hat{g}]=\nabla L(\theta)+b,\quad\|b\|=O\left(\frac{\epsilon}{\sqrt{m}}\right).$ (32)
The bias $b$ decreases with smaller $\epsilon$ and larger $m$.
Assume the loss function $L(\theta)$ is twice continuously differentiable with $\mathcal{L}$-Lipschitz gradients, so the Hessian $H(\theta)=\nabla^{2}L(\theta)$ exists, and $\|\nabla L(\theta^{\prime})-\nabla L(\theta)\|\leq\mathcal{L}\|\theta^{\prime}-\theta\|$ for all $\theta,\theta^{\prime}$. Taylor expansion around $\theta$:
$L(\theta+\epsilon u_{i})=L(\theta)+\epsilon\nabla L(\theta)^{\top}u_{i}+\frac{1}{2}\epsilon^{2}u_{i}^{\top}H(\theta)u_{i}+O(\epsilon^{3}\|u_{i}\|^{3}),$ (33)
where the $O(\epsilon^{3})$ term is bounded by $\mathcal{L}$ and $\|u_{i}\|=\sqrt{d}$. Thus:
@@
$\hat{g}_{i}=\left(\nabla L(\theta)^{\top}u_{i}+\frac{1}{2}\epsilon u_{i}^{\top}H(\theta)u_{i}+O(\epsilon^{2}d^{3/2})\right)u_{i}.$ (34)
Expectation:
@@
$\mathbb{E}[\hat{g}_{i}]=\mathbb{E}[(\nabla L(\theta)^{\top}u_{i})u_{i}]+\frac{1}{2}\epsilon\mathbb{E}[(u_{i}^{\top}H(\theta)u_{i})u_{i}]+O(\epsilon^{2}d^{3/2}).$ (35)
First term:
$\mathbb{E}[(\nabla^{\top}u_{i})u_{i}]=\sum_{k}\nabla_{k}\mathbb{E}[u_{i,k}u_{i}]=\sum_{k}\nabla_{k}e_{k}=\nabla L(\theta),$ (36)
since $\mathbb{E}[u_{i,k}u_{i,j}]=\delta_{kj}$. Second term: $\mathbb{E}[(u_{i}^{\top}Hu_{i})u_{i}]=O(d)$, so $\frac{1}{2}\epsilon\mathbb{E}[(u_{i}^{\top}Hu_{i})u_{i}]=O(\epsilon d)$. Remainder $O(\epsilon^{2}d^{3/2})$, hence:
$\mathbb{E}[\hat{g}_{i}]=\nabla L(\theta)+O(\epsilon d),\quad\mathbb{E}[\hat{g}]=\nabla L(\theta)+O(\epsilon d).$ (37)
Mean squared error:
$\mathbb{E}[\|\hat{g}-\nabla L(\theta)\|^{2}]=\|\mathbb{E}[\hat{g}]-\nabla L(\theta)\|^{2}+\mathbb{E}[\|\hat{g}-\mathbb{E}[\hat{g}]\|^{2}]=O(\epsilon^{2}d^{2})+\frac{1}{m}\mathbb{E}[\|\hat{g}_{1}-\mathbb{E}[\hat{g}_{1}]\|^{2}].$ (38)
Variance term:
$\mathbb{E}[\|\hat{g}_{1}-\mathbb{E}[\hat{g}_{1}]\|^{2}]=\mathbb{E}[\|\hat{g}_{1}\|^{2}]-\|\mathbb{E}[\hat{g}_{1}]\|^{2}.$ (39)
Expected squared norm:
$\|\mathbb{E}[\hat{g}_{1}]\|^{2}=\|\nabla L(\theta)\|^{2}+O(\epsilon^{2}d^{2}).$ (40)
Expansion for leading term:
$\mathbb{E}[\|\hat{g}_{1}\|^{2}]$ $=\mathbb{E}[\|(\nabla^{\top}u_{i}+\frac{1}{2}\epsilon u_{i}^{\top}Hu_{i}+O(\epsilon^{2}))u_{i}\|^{2}]$ (41)
$=d\mathbb{E}[(\nabla^{\top}u_{i})^{2}+\epsilon(\nabla^{\top}u_{i})(u_{i}^{\top}Hu_{i})+O(\epsilon^{2})]=d\|\nabla L(\theta)\|^{2}+O(\epsilon d^{2}).$ (42)
Thus:
$\mathbb{E}[\|\hat{g}_{1}\|^{2}]-\|\mathbb{E}[\hat{g}_{1}]\|^{2}=d\|\nabla L(\theta)\|^{2}-\|\nabla L(\theta)\|^{2}+O(\epsilon d^{2})=(d-1)\|\nabla L(\theta)\|^{2}+O(\epsilon d^{2}).$ (43)
Define $\sigma^{2}=(d-1)\|\nabla L(\theta)\|^{2}+O(\epsilon d^{2})$:
$\mathbb{E}[\|\hat{g}-\nabla L(\theta)\|^{2}]\leq\frac{\sigma^{2}}{m}+O(\epsilon^{2}d^{2})\leq\frac{\sigma^{2}}{m}+O(\epsilon),$ (44)
absorbing $O(\epsilon^{2}d^{2})$ into $O(\epsilon)$ under assumptions.
@@
Assume the loss function $L(\theta)$ is twice differentiable, with Hessian $H(\theta)$. For small perturbations $\epsilon u_{i}$ (where $u_{i}$ is a Rademacher vector, each component independently $\pm 1$ with probability $1/2$), Taylor expansion gives:
$l_{i}=L(\theta_{t}+\epsilon u_{i})=L(\theta_{t})+\epsilon\nabla L(\theta_{t})^{\top}u_{i}+\frac{1}{2}\epsilon^{2}u_{i}^{\top}H(\theta_{t})u_{i}+O(\epsilon^{3}\|u_{i}\|^{3}).$ (45)
Ignore batch stochasticity for simplicity (result holds in expectation); treat $L$ as deterministic. Sample mean: $\bar{l}=\frac{1}{m}\sum_{i=1}^{m}l_{i}$. Sample variance:
@@
where $\operatorname{Var}(l_{i})$ is the variance of $l_{i}$ (for large $m$, $\sigma_{t}^{2}\approx\operatorname{Var}(l_{i})$, but exactly unbiased). Compute $\operatorname{Var}(l_{i})=\mathbb{E}[(l_{i}-\mathbb{E}[l_{i}])^{2}]$. First, expectation:
$\mathbb{E}[l_{i}]=L(\theta_{t})+\mathbb{E}\left[\frac{1}{2}\epsilon^{2}u_{i}^{\top}Hu_{i}\right]+O(\epsilon^{3}d^{3/2})=L(\theta_{t})+\frac{1}{2}\epsilon^{2}\operatorname{Tr}(H)+O(\epsilon^{3}d^{3/2}),$ (48)
since $\mathbb{E}[\nabla^{\top}u_{i}]=0$ ($\mathbb{E}[u_{i}]=0$) and $\mathbb{E}[u_{i}^{\top}Hu_{i}]=\operatorname{Tr}(H)$ ($\mathbb{E}[u_{i,j}u_{i,k}]=\delta_{jk}$). Centered term:
@@
$=\epsilon^{2}\|\nabla L(\theta_{t})\|^{2},$ (53)
since $\mathbb{E}[u_{i,j}^{2}]=1$, $\mathbb{E}[u_{i,j}u_{i,k}]=0$ for $j\neq k$. Second term (higher-order Hessian variance):
@@
since $\operatorname{Var}(u_{i}^{\top}Hu_{i})=O(\|H\|_{F}^{2}d)$ (Frobenius norm bounded by $\mathcal{L}$), and expectation 0. Third term (cross term):
@@
Second part zero ($\mathbb{E}[u_{i}]=0$). First part: $\mathbb{E}[u_{i,l}u_{i}^{\top}Hu_{i}]=\sum_{j,k}H_{jk}\mathbb{E}[u_{i,l}u_{i,j}u_{i,k}]$. For Rademacher, third moments $\mathbb{E}[u_{j}u_{k}u_{l}]=0$ unless all indices equal, yielding $O(\epsilon^{3}d)$. Combining:
$\operatorname{Var}(l_{i})=\epsilon^{2}\|\nabla L(\theta_{t})\|^{2}+O(\epsilon^{3}d)+O(\epsilon^{4}d).$ (56)
@@
$\mathbb{E}[\sigma_{t}^{2}]=\epsilon^{2}\|\nabla L(\theta_{t})\|^{2}+O(\epsilon^{3}d)+O(\epsilon^{4}d/m)\approx\epsilon^{2}\|\nabla L(\theta_{t})\|^{2}+O(\epsilon^{3}),$ (57)
absorbing higher terms for small $\epsilon$ and large $m$. For perturbed point, similar expansion holds at $\theta_{t}+\epsilon_{\text{sam}}$, with $\sigma_{t,\text{pert}}\approx\epsilon\|\nabla L(\theta_{t}+\epsilon_{\text{sam}})\|$, and update normalization follows.
A.4 Equivalence to SAM

### v2

@@ 6 Conclusion
In this paper, we propose ZOSA, a novel zero-order sharpness-aware minimization framework for efficient prompt tuning of large language models in resource-constrained environments. By integrating batched Rademacher perturbations for gradient estimation, adaptive loss-variance scaling for stability, and sharpness-aware mechanisms to target flat minima. Theoretical analysis establishes $O(1/\sqrt{T})$ convergence under smoothness and bounded variance assumptions, with PAC-Bayesian bounds linking sharpness control to enhanced generalization. Empirical evaluations on synthetic high-dimensional functions and zero-order prompt fine-tuning across GLUE benchmarks validate ZOSA’s superiority, showing faster convergence, higher cosine similarity in gradient estimates, and enhanced accuracy/F1 scores compared to adaptive ZO baselines like ZO-AdaMM and evolutionary methods. These results underscore ZOSA’s robustness in noisy, high-dimensional landscapes, making it a practical solution for zero-order LLM adaptation.
@@ A.1 Fast Zeroth-Order Optimization
ZOSA approximates gradients using a finite difference method along random directions. For the current parameters $\theta_{t}\in\mathbb{R}^{d}$ and empirical loss $L(\theta;\mathcal{B}_{t})$, ZOSA estimates the gradient using $m$ random Rademacher directions $u_{i}$, where each component of $u_{i}$ is independently +1+1 or 1-1 with equal probability:
$\hat{g}_{t,i}=\frac{L(\theta_{t}+\epsilon u_{i};\mathcal{B}_{t})-L(\theta_{t};\mathcal{B}_{t})}{\epsilon}u_{i}.$ (30)
This is a one-sided finite difference approximation. The batched estimator is $\hat{g}_{t}=\frac{1}{m}\sum_{i=1}^{m}\hat{g}_{t,i}$.
@@
Using a Taylor expansion around $\theta_{t}$ (assuming twice continuous differentiability),
$L(\theta_{t}+\epsilon u_{i};\mathcal{B}_{t})=L(\theta_{t};\mathcal{B}_{t})+\epsilon\nabla L(\theta_{t};\mathcal{B}_{t})^{\top}u_{i}+\frac{\epsilon^{2}}{2}u_{i}^{\top}H(\theta_{t})u_{i}+O(\epsilon^{3}).$ (31)
Thus,
$\hat{g}_{t,i}=(\nabla L(\theta_{t};\mathcal{B}_{t})^{\top}u_{i})u_{i}+\frac{\epsilon}{2}(u_{i}^{\top}H(\theta_{t})u_{i})u_{i}+O(\epsilon^{2})u_{i}.$ (32)
Taking expectation over $u_{i}\sim\text{Rademacher}^{d}$,
@@
$\mathbb{E}[\hat{g}_{t,i}]=\nabla L(\theta_{t};\mathcal{B}_{t})+\frac{\epsilon}{2}\mathbb{E}\!\left[(u_{i}^{\top}H(\theta_{t})u_{i})u_{i}\right]+O(\epsilon^{2}).$ (33)
The term $(u_{i}^{\top}H(\theta_{t})u_{i})$ is even in $u_{i}$, but multiplied by $u_{i}$ (odd) yields an odd function, so its expectation is exactly zero by symmetry. Therefore, the bias of a single estimator is $O(\epsilon^{2})$. Averaging over $m$ i.i.d. directions gives
$\mathbb{E}[\hat{g}_{t}]=\nabla L(\theta_{t};\mathcal{B}_{t})+b,\quad\|b\|=O(\epsilon^{2}).$ (34)
The bias $b$ is independent of $m$ and decreases quadratically with $\epsilon$.
For the variance term, under the additional $L$-Lipschitz gradient assumption, the same expansion as in the Gaussian case but exploiting Rademacher properties yields (following identical algebraic steps to the classical analysis),
$\mathbb{E}\!\left[\|\hat{g}_{t,i}\|^{2}\right]-\|\mathbb{E}[\hat{g}_{t,i}]\|^{2}=(d-1)\|\nabla L(\theta_{t};\mathcal{B}_{t})\|^{2}+O(\epsilon d^{2}).$ (35)
Since the $u_{i}$ are i.i.d., the variance of the averaged estimator is $1/m$ times the single-sample variance. The mean squared error is bias-squared plus variance:
@@
$\mathbb{E}\!\left[\|\hat{g}_{t}-\nabla L(\theta_{t};\mathcal{B}_{t})\|^{2}\right]$ $=O(\epsilon^{4})+\frac{(d-1)\|\nabla L(\theta_{t};\mathcal{B}_{t})\|^{2}+O(\epsilon d^{2})}{m}$
$=O\!\left(\frac{d}{m}+\epsilon^{2}\right),$ (36)
where the $O(\epsilon^{4})$ bias-squared term is dominated by the higher-order $O(\epsilon^{2})$ remainder. This completes the proof.
@@
This analysis follows and extends the derivation in dang2025fzoo. Assume the loss function $L(\theta;\mathcal{B}_{t})$ is twice differentiable, with Hessian $H(\theta)$. For small perturbations $\epsilon u_{i}$ (where $u_{i}$ is a Rademacher vector, each component independently $\pm 1$ with probability $1/2$), Taylor expansion gives:
$l_{i}=L(\theta_{t}+\epsilon u_{i};\mathcal{B}_{t})=L(\theta_{t};\mathcal{B}_{t})+\epsilon\nabla L(\theta_{t};\mathcal{B}_{t})^{\top}u_{i}+\frac{1}{2}\epsilon^{2}u_{i}^{\top}H(\theta_{t})u_{i}+O(\epsilon^{3}\|u_{i}\|^{3}).$ (37)
Ignore batch stochasticity for simplicity (result holds in expectation); treat $L$ as deterministic.
Sample mean: $\bar{l}=\frac{1}{m}\sum_{i=1}^{m}l_{i}$.
Sample variance:
@@
where $\operatorname{Var}(l_{i})$ is the variance of $l_{i}$ (for large $m$, $\sigma_{t}^{2}\approx\operatorname{Var}(l_{i})$, but exactly unbiased).
Compute $\operatorname{Var}(l_{i})=\mathbb{E}[(l_{i}-\mathbb{E}[l_{i}])^{2}]$.
First, expectation:
$\mathbb{E}[l_{i}]=L(\theta_{t};\mathcal{B}_{t})+\mathbb{E}\left[\frac{1}{2}\epsilon^{2}u_{i}^{\top}Hu_{i}\right]+O(\epsilon^{3}d^{3/2})=L(\theta_{t};\mathcal{B}_{t})+\frac{1}{2}\epsilon^{2}\operatorname{Tr}(H)+O(\epsilon^{3}d^{3/2}),$ (40)
since $\mathbb{E}[\nabla^{\top}u_{i}]=0$ ($\mathbb{E}[u_{i}]=0$) and $\mathbb{E}[u_{i}^{\top}Hu_{i}]=\operatorname{Tr}(H)$ ($\mathbb{E}[u_{i,j}u_{i,k}]=\delta_{jk}$).
Centered term:
@@
$=\epsilon^{2}\|\nabla L(\theta_{t};\mathcal{B}_{t})\|^{2},$ (45)
since $\mathbb{E}[u_{i,j}^{2}]=1$, $\mathbb{E}[u_{i,j}u_{i,k}]=0$ for $j\neq k$.
Second term (higher-order Hessian variance):
@@
since $\operatorname{Var}(u_{i}^{\top}Hu_{i})=O(\|H\|_{F}^{2}d)$ (Frobenius norm bounded by $\mathcal{L}$), and expectation 0.
Third term (cross term):
@@
Second part zero ($\mathbb{E}[u_{i}]=0$). First part: $\mathbb{E}[u_{i,l}u_{i}^{\top}Hu_{i}]=\sum_{j,k}H_{jk}\mathbb{E}[u_{i,l}u_{i,j}u_{i,k}]$. For Rademacher, third moments $\mathbb{E}[u_{j}u_{k}u_{l}]=0$ unless all indices equal, yielding $O(\epsilon^{3}d)$.
Combining:
$\operatorname{Var}(l_{i})=\epsilon^{2}\|\nabla L(\theta_{t};\mathcal{B}_{t})\|^{2}+O(\epsilon^{3}d)+O(\epsilon^{4}d).$ (48)
@@
$\mathbb{E}[\sigma_{t}^{2}]=\epsilon^{2}\|\nabla L(\theta_{t};\mathcal{B}_{t})\|^{2}+O(\epsilon^{3}d).$ (49)
For the perturbed point, a similar expansion holds at $\theta_{t}+\epsilon_{\text{sam}}$, with $\sigma_{t,\text{pert}}\approx\epsilon\|\nabla L(\theta_{t}+\epsilon_{\text{sam}};\mathcal{B}_{t})\|$, and the update normalization follows.
A.4 Proof of Property 3.3
By $L$-smoothness, Taylor expansion gives:
$\Delta l_{i}=\epsilon\langle\nabla L(\theta_{t};B_{t}),u_{i}\rangle+\frac{\epsilon^{2}}{2}u_{i}^{\top}H(\theta_{t};B_{t})u_{i}+R_{i}+n_{i},$ (50)
where $H=\nabla^{2}L(\theta_{t};B_{t})$ with $\|H\|\leq L$, $|R_{i}|\leq\mathcal{O}(\epsilon^{3}d^{3/2}L)$, $\mathrm{Var}[R_{i}]\leq\mathcal{O}(\epsilon^{6}d^{3}L^{2})$, and $\mathrm{Var}[n_{i}]\leq 2V^{2}$.
Then, $\mathbb{E}[\Delta l_{i}]=\mathcal{O}(\epsilon^{2}dL)$ and
$\mathrm{Var}[\Delta l_{i}]=\epsilon^{2}\|\nabla L(\theta_{t};B_{t})\|^{2}+\mathcal{O}(\epsilon^{4}dL^{2})+2V^{2}\leq\epsilon^{2}G^{2}+2V^{2}+\mathcal{O}(\epsilon^{4}dL^{2}).$ (51)
Let $v=\mathrm{Var}[\Delta l_{i}]$. Thus, $\mathbb{E}[\sigma_{t}]\approx\sqrt{v}\approx\epsilon\|\nabla L(\theta_{t};B_{t})\|$ (with bias $\mathcal{O}(\epsilon^{2}\sqrt{d}L)$ assuming $V\ll\epsilon G$ and small $\epsilon$).
$\sigma_{t}^{2}$ is the (biased) empirical variance, equivalent to a U-statistic of order 2 with kernel $h(x,y)=(x-y)^{2}/2$, where $\mathbb{E}[h(X,Y)]=v$.
By Hoeffding decomposition (boucheron2013concentration),
$\mathrm{Var}[\sigma_{t}^{2}]=\mathcal{O}\left(\frac{\zeta_{1}}{m}+\frac{\zeta_{2}}{m^{2}}\right),$ (52)
with $\zeta_{1}=\mathrm{Var}(\mathbb{E}[h(X,Y)\mid X])=\frac{1}{4}(\mathbb{E}[(X-\mu)^{4}]-v^{2})\leq\mathcal{O}(v^{2})$ (by sub-Gaussianity, (vershynin2018high), $\mathbb{E}[|X-\mu|^{4}]\leq Cv^{2}$ for constant $C$), and $\zeta_{2}=\mathrm{Var}(h(X,Y))\leq M^{4}$ where $M=\mathcal{O}(\epsilon G\sqrt{d}+\epsilon^{2}dL+V\sqrt{\log m})$.
Thus,
$\mathrm{Var}[\sigma_{t}^{2}]=\mathcal{O}\left(\frac{v^{2}}{m}+\frac{M^{4}}{m^{2}}\right).$ (53)
By delta method,
$\mathrm{Var}[\sigma_{t}]\leq\mathcal{O}\left(\frac{v}{m}+\frac{M^{4}}{m^{2}v}\right)=\mathcal{O}\left(\frac{\epsilon^{2}G^{2}+V^{2}}{m}+\frac{\epsilon^{2}d^{2}G^{2}}{m^{2}}\right),$ (54)
simplified to the stated bound (log factors absorbed).
For high-probability, Bernstein’s inequality (vershynin2018high) on $\sigma_{t}^{2}$ yields
$|\sigma_{t}^{2}-v|\leq\mathcal{O}\left(v\sqrt{\frac{\log\delta^{-1}}{m}}+M\frac{\log\delta^{-1}}{m}\right),$ (55)
implying
$|\sigma_{t}-\sqrt{v}|$ $\leq\mathcal{O}\left(\sqrt{\frac{v\log\delta^{-1}}{m}}+\frac{M\log\delta^{-1}}{m\sqrt{v}}\right)$ (56)
$=\mathcal{O}\left(\epsilon G\sqrt{\frac{\log\delta^{-1}}{m}}+V\sqrt{\frac{\log\delta^{-1}}{m}}+\frac{\epsilon\sqrt{d}\log\delta^{-1}}{m}\right).$ (57)
A.5 Approximate Equivalence to SAM